The demand for a commodity is given by P(Q + 1) = 231 and the supply is given by P - Q = 11 , find the consumer and produce surplus.

1
Expert's answer
2022-02-28T09:39:06-0500

P(Q+1)=231

PQ+P= 231

"Q= \frac{231-P}{P}"

Supply

P- Q= 11

Q= P-11

At Equilibrium, "Q_s= Q_d"

"\frac{231-P}{P}= P-11"

"P^2-11P= 231-P"

"P^2-10P-231=0"

Solve for the value of P using the quadratic formula

"P= 21"

Q= 21- 11= 10

What producers are willing to sell at.

It occurs when supply intercept Q=0

P- 0= 11

"P_1= 11"

What consumers are willing to pay.

It occurs when Q= 0

P"_2" = 231

Consumer surplus= "\frac{1}{2}\times (P_2-P_0) \times Q"

"\frac{1}{2}\times( 231-21) \times 10" = 1050

Producer Surplus

"\frac{1}{2}\times (P_1-P_0) \times Q"

"\frac{1}{2}\times( 21-11) \times 10" = 50